Write the equation of the ellipse which satisfies the given conditions.

1a. Center (2, 3), horizontal axis 8, vertical axis 4

b. Center (1, -2), horizontal major axis 8, eccentricity

2a. Foci (-2,1) and (4, 1), major axis 10

b. Foci (-3,0) and (-3, 4), minor axis 6

3a. Foci (-2,2) and (4,2), eccentricity

b. Foci (+2, 0), directrices x = +8

4a. Focus (0,0), vertex (5,0), eccentricity 0.5

b. Focus (0,0), vertex (0, 2), eccentricity 0.6

5a. Center (1,3), V(1,-1), and passing through the origin

b. Center (1, 1), V(3, 1), and passing through the origin


1
Expert's answer
2021-07-23T08:34:06-0400

1.

a)

"C(2, 3), 2a=8, 2b=4"


"\dfrac{(x-2)^2}{(8/2)^2}+\dfrac{(y-3)^2}{(4/2)^2}=1"

"\dfrac{(x-2)^2}{16}+\dfrac{(y-3)^2}{4}=1"

b)

"C(1, -2), 2a=8, \varepsilon=\dfrac{c}{a}=\dfrac{\sqrt{a^2-b^2}}{a}=\dfrac{\sqrt{3}}{2}"


"a^2-b^2=\dfrac{3}{4}a^2"

"b=\dfrac{a}{2}=2"

"\dfrac{(x-1)^2}{16}+\dfrac{(y+2)^2}{4}=1"



2.

a)

"F_1(-2,1), F_2(4, 1), 2a=10"

"a=5"

"k=1, h-c=-2, h+c=4"

"C(1, 1)"


"c=\sqrt{a^2-b^2}=3"

"b=4"

"\dfrac{(x-1)^2}{25}+\dfrac{(y-1)^2}{16}=1"


b)

"F_1(-3,0), F_2(-3, 4), 2b=6"

"b=3"

"h=-3, k-c=0, k+c=4"

"C(-3, 2)"


"c=\sqrt{a^2-b^2}=2"

"a=\sqrt{13}"

"\dfrac{(x+3)^2}{9}+\dfrac{(y-2)^2}{13}=1"


3.

a)

"F_1(-2,2), F_2(4, 2), \dfrac{c}{a}=\dfrac{3}{5}"

"k=2,h-c=-2, h+c=4"

"C(1, 2)"


"c=\sqrt{a^2-b^2}=3"

"a=3(\dfrac{3}{5})=5, b=\sqrt{5^2-3^2}=4"


"\dfrac{(x-1)^2}{25}+\dfrac{(y-2)^2}{16}=1"


b)

"F(\pm2,0), x=\pm8"

"k=0,"

"h+ae=2, h-ae=-2"

"h+\dfrac{a}{e}=8, h-\dfrac{a}{e}=-8"

"h=0"

"a=8e, 8e^2=2"

"e=\dfrac{c}{a}=\dfrac{1}{2}, a=4, c=\sqrt{a^2-b^2}=2"

"C(0, 0)"


"b=2\sqrt{3}"


"\dfrac{x^2}{16}+\dfrac{y^2}{12}=1"


4.

a)

"F_1(0, 0)"

"V_1(5,0)"

"e=0.5"


"k=0,"

"h+ae=0, h+a=5"


"h+0.5a=0, h+a=5"

"a=10, h=-5"

or


"h-ae=0, h+a=5"

"h-0.5a=0, h+a=5"

"a=\dfrac{10}{3}, h=\dfrac{5}{3}"




"a=10, h=-5"

"C(-5,0)"


"c=\sqrt{a^2-b^2}=5"

"b=5\sqrt{3}"

"\dfrac{(x+5)^2}{100}+\dfrac{y^2}{75}=1"





"a=\dfrac{10}{3}, h=\dfrac{5}{3}"


"C(\dfrac{5}{3},0)"

"c=\sqrt{a^2-b^2}=\dfrac{5}{3}"

"b=\dfrac{5\sqrt{3}}{3}"

"\dfrac{(x-\dfrac{5}{3})^2}{100/9}+\dfrac{y^2}{75/9}=1"




b)

"F_1(0, 0)"

"V_1(0,2)"

"e=0.6"

"h=0,""k+ae=0, k+a=2"




"k+0.6a=0, k+a=2"

"a=5, k=-3"

or


"k-ae=0, k+a=2"

"k-0.6a=0, k+a=2"

"a=\dfrac{5}{4}, k=\dfrac{3}{4}"





"a=5, k=-3"

"C(0,-3)"

"c=\sqrt{a^2-b^2}=3"

"b=4"

"\dfrac{x^2}{16}+\dfrac{(y+3)^2}{25}=1"





"a=\dfrac{5}{4}, k=\dfrac{3}{4}"



"C(0,\dfrac{3}{4})"

"c=\sqrt{a^2-b^2}=\dfrac{3}{4}"

"b=1"

"\dfrac{x^2}{1}+\dfrac{(y-\dfrac{3}{4})^2}{25/16}=1"






5.

a)

Center "(1, 3)"

Vertices "(1, -1)"


"\dfrac{(x-h)^2}{b^2}+\dfrac{(y-k)^2}{a^2}=1"

Center "(h, k)"

Vertices "(h, k\pm a)"


"h=1, k=3"

"3-a=-1=>a=4"

Passing through the origin


"\dfrac{(0-1)^2}{b^2}+\dfrac{(0-3)^2}{4^2}=1"

"b^2=16\/7"

"\dfrac{(x-1)^2}{16/7}+\dfrac{(y-3)^2}{16}=1"

b)

Center "(1, 1)"

Vertices "(3, 1)"

"\dfrac{(x-h)^2}{a^2}+\dfrac{(y-k)^2}{b^2}=1"

Center "(h, k)"

Vertices "(h\pm a, k)"

"h=1, k=1"

"1+a=3=>a=2"

Passing through the origin


"\dfrac{(0-1)^2}{2^2}+\dfrac{(0-1)^2}{b^2}=1"

"b^2=4\/3"

"\dfrac{(x-1)^2}{4}+\dfrac{(y-1)^2}{4/3}=1"